Betty took out a subsidized student loan of $35,000 at a 4.8% APR, compounded monthly, to pay for her last four semesters of college. If she will begin paying off the loan in 21 months with monthly payments lasting for 20 years, what will be the amount of her monthly payment? 

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Hi Chelsie 

Her amount due = 35,000(1.004)^273 
If her monthly payment is x 
She'll pay = x(1.004)^240 + x(1.004)^239 + ....x(1.004)^1 + x 
= x[(1.004)^241 - 1]/(1.004-1) = 
Her payment = Amount due 
so 
x[(1.004)^241 - 1]/(1.004-1) = 35000(1.004)^273 
404x = 35000(1.004)^273 = 104081
x = 104081/404 = $257.6

Monthly payment = $257.6
 
Regards
Arun (askIITians forum expert)